OpenStudy (anonymous):

Find the derivative of :

OpenStudy (anonymous):

\[\huge (\frac{ 1 }{ 3 }) ^{4x} -2e^{sinx} \]

OpenStudy (anonymous):

the derivative of \(b^x\) is \(b^x\ln(b)\) that should help for the first one

OpenStudy (anonymous):

that, plus chain rule (for both)

OpenStudy (anonymous):

\[\large y'=(4\ln \frac{ 1 }{ 3 }) (\frac{ 1 }{ 3 })^{4x}-2e^{sinx}(cosx)\] @satellite73

OpenStudy (anonymous):

That's the answer at the back of my book but where did the 4 before ln comefrom ?